Accelerator Update for January 18, 2008 - January 21, 2008
Friday January 18
The day shift began with Operations monitoring the Tevatron (TeV) store 5852 (58.04E30) and the Recycler stash (324.7E10), with the Antiproton Source not stacking due to problems with a power supply (A:IB), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.
Operations reported at 8:11 AM that they were running a NuMI only timeline during the Pbar A:IB investigations.
At 8:45 AM, a Pbar power supply expert reported that he had found a ground fault on A:IB. At 9:17 AM, he reported that the ground fault was in the tunnel.
Operations began preparing Pbar’s Prevault area for an access at 9:22 AM.
At 9:24 AM, a Linac RF station (LRF2) tripped off. It reset.
At 10:22 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.
The Pbar power supply expert reported at 11:07 AM that he found the Accumulator device (A:IB) ground fault; it was ground water dripping on an insulator holding the buss. Pbar came out of access at 11:29 AM.
At 11:48 AM, MIRF8 tripped off. It reset.
The expert had A:IB back online by 11:56 AM.
At 12:11 PM, MIRF18 tripped off and would not reset. An expert investigated and reported that the station needed a new bias supply.
An RF expert had MIRF18 back on by 12:56 PM.
At 1:03 PM, a Pbar expert reported that DRF1-7 was slowly failing.
Operations terminated store 5852 at 1:34 PM.
At 1:39 PM, Operations reported that LRF2 had tripped off three times. A Linac expert investigated and reported that one of the switch tubes was sparking. They were bringing the station back up slowly.
At 1:44 PM, other Pbar experts investigated problems with DRF1-7.
Operations put Pbar in standby at 1:59 PM; it will stay in standby until LRF2 comes back online.
Pbar resumed stacking at 3:30 PM.
Operations put the TeV into shot setup at 3:33 PM.
Operations established store 5853 at 5:53 PM with an initial luminosity of 180.84E30.
At 7:57 PM, LRF2 tripped off for the first of three times in about 90 minutes. It reset.
At 11:10 PM, LRF5 tripped off for the first of four times in five minutes. It reset.
Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.
Saturday January 19
The midnight shift began with Operations monitoring store 5853 (88.75E30) and the stash (157.1E10), with Pbar stacking (17.2E10), and with NuMI and BNB taking beam.
At 4:32 AM, Operations reported that the Pelletron started tripping off. Operators contacted an expert.
A Pelletron expert reported at 7:38 AM that he got the Pelletron up and running, but the regulation would not hold. He, mechanical technicians, and operators began preparing the MI-31 area for access.
Operations reported at 8:02 AM that the Recycler had stashed once during the midnight shift.
At 8:49 AM, BRF1 tripped off for the first of three times in three hours.
At 11:12 AM, the Booster beam permit dropped on a vacuum alarm. A vacuum valve indicated that it was moving. Operators investigated and everything seemed to be normal. They contacted experts. Operations turned off the Booster beam.
At 11:47 AM, LRF 2 tripped off for the first of two times in five minutes. It reset.
Operations began at 1:09 PM preparing MI for an access to search for a ground fault.
Operations reported at 1:28 PM that a Booster expert was able to circulate beam around the Booster.
MI experts reported at 2:45 PM that there was still a ground fault, but it wasn’t as bad as they first thought. The experts thought it would hold until after the next shot to the TeV.
Operations resumed sending beam to BNB around 3:40 PM.
Pbar resumed stacking at 3:45 PM. Operations reported that NuMI was being held off due to a problem with a low oil level on a transformer (E:V108). Operations called in an expert who had E:V108 repaired by 5:16 PM.
Operations terminated store 5853 at 5:31 PM.
Operation put the TeV into shot setup at 5:41 PM.
Operations established store 5854 at 7:36 PM with an initial luminosity of 105.3E30.
Operations prepared MI for an access at 7:45 PM.
A MI expert reported at 9:36 PM that they found a LCW leak on a quadrupole (Q631) was the cause of the ground fault. The experts temporarily rerouted the leak.
MI came out of access at 11:07 PM.
Sunday January 20
The midnight shift began with Operations monitoring store 5854 and the stash, with Pbar, NuMI, and BNB off due to the MI access.
Operations established beam to BNB at 12:09 AM and to NuMI at 12:22 AM.
Operations established beam to Pbar at 1:16 AM. The delay was due to a problem with Pbar Debuncher kicker (D:EKIK).
Operations terminated store 5854 a 5:26 AM.
Operations prepared D0 and CDF for access at 6:11 AM.
Operators search and secured MI-31 at 9:14 AM.
Operations reported at 11:28 AM that temperatures in the Booster West Tower offices had dropped. Two water pipes had frozen and burst. FESS valved off the piping involved and installed portable heaters.
D0 and CDF were out of access by 5:10 PM.
Operations reported at 7:57 PM that the Recycler had stashed twice during the last twelve hours.
At 9:09 PM, Operations reported that they were replacing two TeV corrector magnet power supplies (HC48 and VB16).
Operations put the TeV through a dry squeeze at 11:21 PM.
Monday January 21
The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and BNB taking beam.
Operations put the TeV through a wet squeeze at 12:31 AM.
Operations put the TeV into shot setup at 12:59 AM.
Operations established store 5857 at 2:56 AM with an initial luminosity of 149.07E30.
At 3:22 AM, all the Booster correctors began working erratically. Operators investigated and found water dripping on the power supplies. Operations turned off Booster, breakered off power supplies, and contacted experts.
At 4:53 AM, Operations reported that another water pipe in the Booster Tower had frozen and cracked in an upstairs office area and dripped through the floor to the power supplies below. Operators called in many experts.
Operations resumed sending beam to NuMI, BNB, and Pbar at 7:52 AM.
The Plans for Monday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 01/24/2008 email Fermilab|
|Security, Privacy, Legal|